OpenStudy (anonymous):

Find the sum of the infinite geometric series, if it exists. 4-1+1/4-1/16+... A. - 1 B. 3 C. 16/5 D. does not exist Solve the problem. Round to the nearest dollar if needed. Looking ahead to retirement, you sign up for automatic savings in a fixed-income 401K plan that pays 5% per year compounded annually. You plan to invest $3500 at the end of each year for the next 15 years. How much will your account have in it at the end of 15 years? A. $77,295 B. $75,525 C. $76,823 D. $73,982 Find the term indicated in the expansion. (x - 3y)^11; 8th term A. -721,710x^7y^4 B. -721,710x^4y^7 C. 240,570x^7y^4 D. 240,570x^4y^8

OpenStudy (raden):

use the formula : S = a/(1-r) in this case, a=4, r=-1/4

OpenStudy (anonymous):

@RadEn what about the other two?